OpenStudy (anonymous):

A map has dimensions of 9 in. by 15 in. You want to reproduce the map so that it will fit on a 4 in. by 6 in. index card. What are the dimensions of the largest possible complete map that you can fit on the index card? a. 3.2 in. by 6 in. c. 3.9 in. by 6 in. b. 3.6 in. by 6 in d. 4 in. by 6 in.

OpenStudy (anonymous):

wouldn't it be D because you want to reproduce it on that

OpenStudy (xguardians):

Check to see which ones equal the same thing. The ratio for the original is 9/15, so find which others are equal, but close to it. 9/15 = .6 (Original) -------------- 3.2/6 = .53 (A) 3.9 / 6 = .65 (b) 3.6 / 6 = .6 (C) <---- DING MATCH! 4 / 6 = .66 (D) Since they have the same ratio, the answer is C.